# A Unified Implementation of Quasi-Monte Carlo Generators, Randomization Routines, and Fast Kernel Methods

### v1

@@ 1.2. Kernel computations
Low-discrepancy (LD) point sets can also been used to accelerated many kernel computations. Let $\{{\boldsymbol{x}}_{i}\}_{i=0}^{n-1}$ be a point set, $K:[0,1]^{d}\times[0,1]^{d}\to{\mathbb{R}}$ be a symmetric positive definite (SPD) kernel, and ${\mathsf{K}}=\left(K({\boldsymbol{x}}_{i},{\boldsymbol{x}}_{k})\right)_{i,k=0}^{n-1}$ be the $n\times n$ SPD Gram matrix. Two motivating kernel computations are described below
Discrepancy Computation:
@@
The error of the Monte Carlo approximation (1) can be bounded by the product two terms: a measure of variation of the function $f$ and a measure of the discrepancy of the point set $\{{\boldsymbol{x}}_{i}\}_{i=1}^{n-1}$. The most well known bound of this type is the Koksma-Hlawka inequality (Hickernell, 1998; Dick et al., 2013; Hickernell, 1999; Niederreiter, 1992). The discrepancy is frequently computed when designing and evaluating low-discrepancy point sets, see (Rusch et al., 2024) for a newer idea in this area where good point sets are generated using a neural networks trained with a discrepancy-based loss function.
@@
Let us consider the more general setting of approximating the mean $\mu$ in (1) by a weighted cubature rule $\sum_{i=0}^{n-1}\omega_{i}f({\boldsymbol{x}}_{i})$. If $f$ is assumed to lie in the RKHS $H$, the discrepancy in the error bound of such a cubature rule takes the form
$\int\int K({\boldsymbol{u}},{\boldsymbol{v}})\mathrm{d}{\boldsymbol{u}}\mathrm{d}{\boldsymbol{v}}-2\sum_{i=0}^{n-1}\omega_{i}\int K({\boldsymbol{u}},{\boldsymbol{x}}_{i})\mathrm{d}{\boldsymbol{u}}+\sum_{i,k=0}^{n-1}\omega_{i}\omega_{k}K({\boldsymbol{x}}_{i},{\boldsymbol{x}}_{k})$
@@
following (Hickernell, 1998). Here the integrals are understood to be over the unit cube $[0,1]^{d}$. Evaluating this discrepancy above requires computing ${\mathsf{K}}{\boldsymbol{\omega}}$ where ${\boldsymbol{\omega}}=\{\omega_{i}\}_{i=0}^{n-1}$. Moreover, the discrepancy is minimized for a given point set by setting the weights to ${\boldsymbol{\omega}}^{\star}={\mathsf{K}}^{-1}{\boldsymbol{\kappa}}$ where $\kappa_{i}=\int K({\boldsymbol{u}},{\boldsymbol{x}}_{i})\mathrm{d}{\boldsymbol{u}}$ for $0\leq i<n$.
Kernel Interpolation:
@@
Suppose we would like to approximate $f:[0,1)^{d}\to{\mathbb{R}}$ given observations ${\boldsymbol{y}}=\{y_{i}\}_{i=0}^{n-1}$ of $f$ at $\{{\boldsymbol{x}}_{i}\}_{i=0}^{n-1}$ satisfying $y_{i}=f(x_{i})$. Then a kernel interpolant approximates $f$ by ${\widehat{f}}\in H$ where
@@
and ${\boldsymbol{\omega}}={\mathsf{K}}^{-1}{\boldsymbol{y}}$. The above kernel interpolant may be reinterpreted as the posterior mean in Gaussian process regression, see (Williams and Rasmussen, 2006). Fitting a Gaussian process regression model often includes optimizing a kernel’s hyperparameters, which also may be done by computing ${\mathsf{K}}{\boldsymbol{\omega}}$ and ${\mathsf{K}}^{-1}{\boldsymbol{y}}$.
Underlying these problems, and many others, is the requirement to compute the matrix-vector product ${\mathsf{K}}{\boldsymbol{y}}$ or solve the linear system ${\mathsf{K}}\boldsymbol{\omega}={\boldsymbol{y}}$ for some length $n$ vector ${\boldsymbol{y}}$. The standard cost of these computations are ${\mathcal{O}}(n^{2})$ and ${\mathcal{O}}(n^{3})$ respectively. One method to reduce these high costs is to induce structure into ${\mathsf{K}}$. Structure-inducing methods pairing lattices and digital nets with special kernels were proposed in (Zeng et al., 2009) and shown to reduce the cost of both computations to ${\mathcal{O}}(n\log n)$. Specifically, when a lattice point set $\{{\boldsymbol{x}}_{i}\}_{i=0}^{n-1}$ in linear order is paired with a shift-invariant kernel, ${\mathsf{K}}$ becomes circulant and computations can be performed performed using the fast Fourier transforms (FFT) and inverse FFT (IFFT). Similarly, when a base $2$ digital net $\{{\boldsymbol{x}}_{i}\}_{i=0}^{n-1}$ is paired with a digitally-shift-invariant kernel, ${\mathsf{K}}$ becomes nested-block-Toeplitz and these computations can be performed using the Fast Walsh-Hadamard Transforms (FWHT) (Fino and Algazi, 1976). Such methods have recently been used to accelerate Bayesian cubature (Jagadeeswaran and Hickernell, 2019, 2022; Rathinavel, 2019) and solving PDEs with random coefficients (Kaarnioja et al., 2022, 2023).
We provide novel methods which exploit past solutions to ${\mathsf{K}}{\boldsymbol{y}}$ or ${\mathsf{K}}^{-1}{\boldsymbol{y}}$ when doubling the sample size from the LD sequence. This requires lattices are generated in natural order so we do not need to reorder outputs after doubling the sample size. We show that the same speedups can be achieved for lattices in natural order using bit-reversed FFT and bit-reversed IFFT algorithms. These bit-reversed algorithms simply skip the initial and final bit-reversal steps in the decimation-in-time FFT and IFFT respectively. We provide implementations of the bit-reversed FFT, bit-reversed IFFT, and FWHT.
Shift-invariant kernels of arbitrary smoothness are well known and can be computed based on the Bernoulli polynomials (Kaarnioja et al., 2022, 2023; Cools et al., 2021, 2020; Sloan and Woźniakowski, 2001; Kuo et al., 2004). Until now, digitally-shift-invariant kernels were only known for order $1$ smoothness (Dick and Pillichshammer, 2005). We present higher-order digitally-shift-invariant kernels whose corresponding RKHSs contain functions of arbitrary smoothness. We derive closed forms of low order kernels (up to order $4$ smoothness) in base $b=2$, with the order $4$ kernel not appearing elsewhere in the literature. These kernels are also of interest for computing the worst-case error of higher-order polynomial-lattice rules (Baldeaux et al., 2012), although we do not explore this connection here.
1.3. Outline
The remainder of this article is organized as follows. First, we describe our implementation in section 2. Section 3 discusses notation. Section 4 details rank $1$ lattices, digital sequences, and Halton sequences along with their randomization routines. Section 5 describes accelerated kernel computation with LD sequences, matching kernels, and fast transform algorithms. Section 6 provides numerical experiments. Finally, section 7 gives a brief conclusion and outlines directions for future work.
2. Implementation
We have implemented this work into QMCPy [^10] https://qmcpy.readthedocs.io version 1.6, installable with pip install -U qmcpy. The architecture and philosophy underlying QMCPy are described in (Choi et al., 2022b, a). To follow along with code snippets presented in this article, please import QMCPy and NumPy (Oliphant, 2006) via
@@ 2. Implementation
These code snippets and code producing the plots in this article are available in notebook form [^11] https://github.com/QMCSoftware/QMCSoftware/blob/master/demos/ld_randomizations_and_higher_order_nets.ipynb.
Most of the QMCPy objects and methods presented here are wrappers around the lower level QMCToolsCL [^12] https://qmcsoftware.github.io/QMCToolsCL/ package. The Python functions in QMCToolsCL often call lower level C implementations for improved performance. QMCToolsCL interfaces for other languages which support C extensions is a direction of future work.
3. Notation
Bold symbols denote vectors which are assumed to be column vectors e.g. ${\boldsymbol{x}}\in[0,1]^{d}$. Capital letters in serif font denote matrices e.g. ${\mathsf{K}}\in{\mathbb{R}}^{n\times n}$. Lower case letters in serif font denote digits in a base $b$ expansion e.g. $i=\sum_{t=0}^{m-1}{\mathsf{i}}_{t}b^{t}$ for $0\leq i<2^{m}$. This may be combined with bold notation when denoting the vector of base $b$ digits e.g. ${\boldsymbol{\mathsf{i}}}=({\mathsf{i}}_{0},{\mathsf{i}}_{1},\dots,{\mathsf{i}}_{m-1})^{\intercal}$. Modulo is always taken component-wise e.g. ${\boldsymbol{x}}\mod 1=(x_{1}\mod 1,\dots,x_{d}\mod 1)^{\intercal}$. Permutations may be specified by vectors e.g. $\pi:\{0,1,2\}\to\{0,1,2\}$ with $\pi(0)=2$, $\pi(1)=0$, and $\pi(2)=1$ may be denoted by $\pi=(2,0,1)$.
4. Low-discrepancy sequences and randomization routines
@@
For ${\boldsymbol{\mathsf{i}}}={\boldsymbol{D}}_{m}(i)$ we use the notation
$F_{m}({\boldsymbol{\mathsf{i}}})=\sum_{t=1}^{m}{\mathsf{i}}_{t-1}b^{-t}$
to flip the digits of $i$ about the decimal and convert back to a floating point number in $[0,1)$. Finally, let
(3) $v(i)=\sum_{t=1}^{\infty}{\mathsf{i}}_{t-1}b^{-t}$
so that $v(i)=F_{m}({\boldsymbol{D}}_{m}(i))$ when $i<b^{m}$, which is always the case in this article. The Van der Corput sequence in base $b$ is $(v(i))_{i\geq 0}$.
@@
$x$$z$shifted digital net
$x$$z$permuted digital net
$x$$z$NUS digital net
Figure 2. Low-discrepancy randomization routines in dimension $d=1$. Each vertical line is a unit interval with 00 at the bottom and $1$ at the top. A given interval is partitioned at the horizontal ticks extending to the right, and then rearranged following the arrows to create the partition of the right interval as shown by the ticks extending to the left. For the shifted lattice and digitally shifted digital net, when the arrow hits a horizontal gray bar it is wrapped around to the next gray bar below. See for example the dotted line in the shifted digital net. The lattice shift is $\Delta=0.227$. All digital nets use base $b=3$ and $t_{\mathrm{max}}=2$ digits of precision. The digital shift is ${\boldsymbol{\Delta}}=(2,1)^{\intercal}$. Dropping the $j$ subscript for dimension, the digital permutations in the third panel are $\pi_{1}=(2,0,1)$ and $\pi_{2}=(2,1,0)$. Notice $\pi_{1}$ is equivalent to a digital shift by $2$, but $\pi_{2}$ cannot be written as a digital shift. The nested uniform scramble (NUS) has digital permutations $\pi=(2,0,1)$, $\pi_{0}=(2,1,0)$, $\pi_{1}=(0,1,2)$, and $\pi_{2}=(1,2,0)$. Notice the permuted digital net in third panel has permutations depending only the current digit in the base $b$ expansion. In contrast, the full NUS scrambling in the fourth panel has permutations which depend on all previous digits in the base $b$ expansion.
4.1. Rank 1 lattices
@@
${\boldsymbol{z}}_{i}=v(i){\boldsymbol{g}}\mod 1,\qquad i\geq 0.$
If $n=b^{m}$ for some $m\in{\mathbb{N}}_{0}$, then the lattice point set $\{{\boldsymbol{z}}_{i}\}_{i=0}^{n-1}\in[0,1)^{n\times d}$ is equivalent to $\{i/n{\boldsymbol{g}}\mod 1\}_{i=0}^{n-1}$ where we say the former is in natural order while the latter is in linear order.
Shifted lattice:
@@
where ${\boldsymbol{z}}_{i}$ are lattice point set. The shift operation mod 1 is visualized in fig. 2. Randomized lattices use ${\boldsymbol{\Delta}}\sim{\mathcal{U}}[0,1]^{d}$. In the following code snippet we generate $R$ independently shifted lattices with shifts ${\boldsymbol{\Delta}}_{1},\dots,{\boldsymbol{\Delta}}_{R}\overset{\mathrm{IID}}{\sim}{\mathcal{U}}[0,1]^{d}$.
@@
randomize = "shift", # for unrandomized lattice set randomize = False
@@
order = "natural", # also supports "linear" and "Gray" for Gray code
@@
generating_vector = "kuo.lattice-33002-1024-1048576.9125.txt")
x = lattice(2**16) # a numpy.ndarray with shape 16 x 65536 x 52
Here we have used a generating vector from (Cools et al., 2006) which is stored in a standardized format in the LDData repository. Other generating vectors from LDData may be used by passing in a file name from https://github.com/QMCSoftware/LDData/tree/main/lattice/ or by passing an explicit array.
4.2. Digital nets
Consider a fixed prime base $b$ and generating matrices ${\mathsf{C}}_{1},\dots\mathsf{C}_{d}\in\{0,...,b-1\}^{t_{\mathrm{max}}\times m}$ where $m\in{\mathbb{N}}$ is fixed and $b$ is a given prime base. One may relax the assumption that $b$ is prime, but we do not consider that here or in the implementation. The first $n=b^{m}$ points of a digital sequence form a digital net $\{{\boldsymbol{z}}_{i}\}_{i=1}^{b^{m}-1}\in[0,1)^{b^{m}\times d}$ in natural order. For $1\leq j\leq d$ and $0\leq i<b^{m}$ with digits ${\boldsymbol{\mathsf{i}}}={\boldsymbol{D}}_{m}(i)$,
${\boldsymbol{\mathsf{z}}}_{ij}={\mathsf{C}}_{j}{\boldsymbol{\mathsf{i}}}\mod b\qquad\text{and}\qquad z_{ij}=F_{t_{\mathrm{max}}}({\boldsymbol{\mathsf{z}}}_{ij})$
where ${\boldsymbol{\mathsf{z}}}_{ij}$ is the base $b$ expansion vector of $z_{ij}$.
Digitally shifted digital net:

### v2

@@ 3.2 Fast Kernel computations
Randomized LD point sets have also been used to accelerated kernel computations. Let $\{{\boldsymbol{x}}_{i}\}_{i=0}^{n-1}$ be a point set, $K:[0,1]^{d}\times[0,1]^{d}\to{\mathbb{R}}$ be a symmetric positive definite (SPD) kernel, and ${\mathsf{K}}=\left(K({\boldsymbol{x}}_{i},{\boldsymbol{x}}_{i^{\prime}})\right)_{i,i^{\prime}=0}^{n-1}$ be the $n\times n$ SPD Gram matrix of kernel evaluations at all pairs of points. Two motivating kernel computations are described below
Discrepancy Computation
@@
The error of the approximation (1) can be bounded by the product of two terms: a measure of variation of the function $f$ and a measure of the discrepancy of the point set $\{{\boldsymbol{x}}_{i}\}_{i=0}^{n-1}$. The most well known bound of this type is the Koksma–Hlawka inequality [Hickernell, 1998a, Dick et al., 2013, Hickernell, 1999, Niederreiter, 1992]. The discrepancy is frequently computed when designing and evaluating low-discrepancy point sets, see [Rusch et al., 2024] for a newer idea in this area where good point sets are generated using neural networks trained with a discrepancy-based loss function.
@@
Let us consider the more general setting of approximating the mean $\mu$ in (1) by a weighted cubature rule $\sum_{i=0}^{n-1}\omega_{i}f({\boldsymbol{x}}_{i})$. If $f$ is assumed to lie in the RKHS $H$ with kernel $K$, the discrepancy in the error bound of such a cubature rule takes the form
$\int_{[0,1]^{d}}\int_{[0,1]^{d}}K({\boldsymbol{u}},{\boldsymbol{v}})\mathrm{d}{\boldsymbol{u}}\mathrm{d}{\boldsymbol{v}}-2\sum_{i=0}^{n-1}\omega_{i}\int_{[0,1]^{d}}K({\boldsymbol{u}},{\boldsymbol{x}}_{i})\mathrm{d}{\boldsymbol{u}}+\sum_{i,i^{\prime}=0}^{n-1}\omega_{i}\omega_{i^{\prime}}K({\boldsymbol{x}}_{i},{\boldsymbol{x}}_{i^{\prime}})$
@@
following [Hickernell, 1998a]. Evaluating the discrepancy above requires computing ${\mathsf{K}}{\boldsymbol{\omega}}$ where ${\boldsymbol{\omega}}=\{\omega_{i}\}_{i=0}^{n-1}$. Moreover, this discrepancy is minimized for a given point set by setting the weights to ${\boldsymbol{\omega}}^{\star}={\mathsf{K}}^{-1}{\boldsymbol{\kappa}}$ where $\kappa_{i}=\int_{[0,1]^{d}}K({\boldsymbol{u}},{\boldsymbol{x}}_{i})\mathrm{d}{\boldsymbol{u}}$ for $0\leq i<n$.
Kernel Interpolation
@@
Suppose we would like to approximate $f:[0,1)^{d}\to{\mathbb{R}}$ given observations ${\boldsymbol{y}}=\{y_{i}\}_{i=0}^{n-1}$ of $f$ at $\{{\boldsymbol{x}}_{i}\}_{i=0}^{n-1}$ satisfying $y_{i}=f(x_{i})$ for $0\leq i<n$. Then a kernel interpolant approximates $f$ by ${\widehat{f}}\in H$ where
@@
and ${\boldsymbol{\omega}}={\mathsf{K}}^{-1}{\boldsymbol{y}}$. The above kernel interpolant may be reinterpreted as the posterior mean of a Gaussian process regression model, see [Williams and Rasmussen, 2006]. Fitting such a Gaussian process often includes optimizing a kernel’s hyperparameters, which may also be done by computing ${\mathsf{K}}{\boldsymbol{\omega}}$ and ${\mathsf{K}}^{-1}{\boldsymbol{y}}$.
Underlying these problems, and many others, is the requirement to compute the matrix-vector product ${\mathsf{K}}{\boldsymbol{y}}$ or solve the linear system ${\mathsf{K}}^{-1}{\boldsymbol{y}}$. The standard cost of these computations are ${\mathcal{O}}(n^{2})$ and ${\mathcal{O}}(n^{3})$ respectively. One method to reduce these high costs is to induce structure into $\mathsf{K}$}. Zeng et al. [2009, 2006] proposed structure-inducing methods which use certain kernels $K$ and points $\{{\boldsymbol{x}}_{i}\}_{i=0}^{n-1}$ which enable both computations to be done at only ${\mathcal{O}}(n\log n)$ cost. Two such pairings exist:
  1. 1.
    When a rank-1 lattice point set $\{{\boldsymbol{x}}_{i}\}_{i=0}^{n-1}$ in linear order is paired with a SI kernel, $\mathsf{K}$} is circulant and thus diagonalizable by the FFT and IFFT. Using lattices in the extensible radical inverse order requires using the bit-reversed fast transforms FFTBR and IFFTBR.
  2. 2.
    When a base $2$ digital net $\{{\boldsymbol{x}}_{i}\}_{i=0}^{n-1}$ in radical inverse order is paired with a DSI kernel, $\mathsf{K}$} becomes RSBT (recursive symmetric block Toeplitz) and thus diagonalizable by the FWHT.
4 Implementation
The features discussed in this article are available in QMCPy version 2.0 or later, which may be readily installed using the command pip install -U qmcpy. To follow along with the code snippets in this paper, first
Listing 1: Import the QMCPy and NumPy Python packages to be used in examples throughout this paper.
@@ 4 Implementation
All codes and plots are reproducible in a public notebook [^14] https://qmcsoftware.github.io/QMCSoftware/demos/talk_paper_demos/ACMTOMS_Sorokin_2025/acm_toms_sorokin_2025/. Most of the QMCPy functions presented in this paper are wrappers around efficient C implementations in our QMCToolsCL [^15] https://qmcsoftware.github.io/QMCToolsCL/ package. These low level routines could be used to rapidly build similar interfaces in other programming languages which support C extensions.
5 Randomized low-discrepancy sequences
@@
For ${\boldsymbol{\mathsf{i}}}={\boldsymbol{D}}_{m}(i)$, we denote the radical inverse of $i$ by
$F_{m}({\boldsymbol{\mathsf{i}}})=\sum_{t=1}^{m}{\mathsf{i}}_{t-1}b^{-t}\in[0,1)$
Finally, let
$v(i)=\sum_{t=1}^{\infty}{\mathsf{i}}_{t-1}b^{-t}$ (3)
so that $v(i)=F_{m}({\boldsymbol{D}}_{m}(i))$ when $i<b^{m}$, which is always the case in this article. The van der Corput sequence in base $b$ is $\{v(i)\}_{i\geq 0}$.
@@
xxzzdigitally shifted DN
xxzzpermuted DN
xxzzNUS DN
Figure 2: Low-discrepancy randomization routines in dimension $d=1$. Each vertical line is a unit interval with 0 at the bottom and 11 at the top. A given interval is partitioned at the horizontal ticks extending to the right, and then rearranged following the arrows to create the partition of the right interval as shown by the ticks extending to the left. For the shifted rank-1 lattice and digitally shifted digital net, when the arrow hits a horizontal gray bar it is wrapped around to the next gray bar below. See for example the dotted line in the digitally shifted digital net. The lattice shift is $\Delta=0.227$. All digital nets (DNs) use base $b=3$ and $t_{\mathrm{max}}=2$ digits of precision. The digital shift is ${\boldsymbol{\Delta}}=(2,1)^{\intercal}$. Dropping the $j$ subscript for dimension, the digital permutations in the third panel are $\pi_{1}=(2,0,1)$ and $\pi_{2}=(2,1,0)$. Notice $\pi_{1}$ is equivalent to a digital shift by 22, but $\pi_{2}$ cannot be written as a digital shift. The nested uniform scramble (NUS) has digital permutations $\pi=(2,0,1)$, $\pi_{0}=(2,1,0)$, $\pi_{1}=(0,1,2)$, and $\pi_{2}=(1,2,0)$. Notice the permuted digital net in the third panel has permutations depending only on the current digit in the base $b$ expansion. In contrast, the full NUS scrambling in the fourth panel has permutations which depend on all previous digits in the base $b$ expansion.
5.1 Rank-1 lattices
@@
${\boldsymbol{z}}_{i}=v(i){\boldsymbol{g}}\mod 1,\qquad i\geq 0.$ (4)
If $n=b^{m}$ for some $m\in{\mathbb{N}}_{0}$, then the lattice point set $\{{\boldsymbol{z}}_{i}\}_{i=0}^{n-1}\in[0,1)^{n\times d}$ is equivalent to $\{{\boldsymbol{g}}i/n\mod 1\}_{i=0}^{n-1}$ where we say the former is in radical inverse order while the latter is in linear order. When $n$ is not of the form $b^{m}$, then linear and radical inverse order will not generate the same lattice point set.
Shifted lattice
@@
where ${\boldsymbol{z}}_{i}$ denotes the unrandomized lattice point in (4). The shift operation mod 1 is visualized in Figure 2. Randomized lattices use ${\boldsymbol{\Delta}}\sim{\mathcal{U}}[0,1]^{d}$. In the following code snippet we generate $R$ independently shifted lattices with shifts ${\boldsymbol{\Delta}}_{1},\dots,{\boldsymbol{\Delta}}_{R}\overset{\mathrm{IID}}{\sim}{\mathcal{U}}[0,1]^{d}$.
Listing 2: Generate randomized lattices.
@@
randomize = "shift", # for unrandomized set randomize = None
@@
order = "radical inverse", # also supports "linear"
@@
generating_vector = "mps.exod2_base2_m20_CKN.txt")
x = lattice(2**16) # a numpy.ndarray with shape (16, 65536, 52)
Here we have used a generating vector from [Cools et al., 2006] which is stored in a standardized format in the LDData repository. Other generating vectors from LDData may be used by passing in a file name from https://github.com/QMCSoftware/LDData/tree/main/lattice/ or by passing in an explicit array.
5.2 Digital nets
Consider a fixed prime base $b$ and generating matrices ${\mathsf{C}}_{1},\dots\mathsf{C}_{d}\in\{0,...,b-1\}^{t_{\mathrm{max}}\times m}$ where $t_{\mathrm{max}},m\in{\mathbb{N}}$ are fixed and $b$ is a given prime base. One may relax the assumption that $b$ is prime, but we do not consider that here or in the implementation. The first $n=b^{m}$ points of a digital sequence form a digital net $\{{\boldsymbol{z}}_{i}\}_{i=0}^{b^{m}-1}\in[0,1)^{b^{m}\times d}$ where, in radical inverse order, for $1\leq j\leq d$ and $0\leq i<b^{m}$ with base $b$ digit vector ${\boldsymbol{\mathsf{i}}}={\boldsymbol{D}}_{m}(i)$,
$z_{ij}=F_{t_{\mathrm{max}}}({\boldsymbol{\mathsf{z}}}_{ij})\qquad\text{with}\qquad{\boldsymbol{\mathsf{z}}}_{ij}={\mathsf{C}}_{j}{\boldsymbol{\mathsf{i}}}\mod b.$
Digitally-shifted digital net